# Fixed-Point RNNs: Interpolating from Diagonal to Dense

### v2

@@ E.2 Long-Range State-Tracking
((a)) State Tracking on $A_{5}$ FPMamba
((b)) State Tracking on $S_{5}$ FPMamba
((c)) State Tracking on $A_{5}$ FPMamba2
((d)) State Tracking on $S_{5}$ FPMamba2
((e)) State Tracking on $A_{5}$ DeltaProduct
((f)) State Tracking on $S_{5}$ DeltaProduct
Figure 9: The state-tracking experiment for train sequence length $128$ and evaluation sequence length $[2,512]$. We omit the results of DPLR mixer due to poor performance. The figure presents (a, b) the results for FPMamba with Householder (H) and Kronecker (K) mixer, (c, d) the results for FPMamba2 with Householder (H) and Kronecker (K) mixer, and (e, f) the DeltaProduct method (Siems et al., 2025) for Householder mixers.
@@
In this section, we investigate the ability of our proposed method in doing state tracking on longer sequences. Specifically, we will use the $A_{5}$ and $S_{5}$ datasets and train on sequence length 128, while evaluating for sequence lengths in the range $[2,512]$. We also implement the proposed Fixed-Point framework on Mamba2 (Dao & Gu, 2024), and we compared our method to DeltaProduct (Siems et al., 2025). In Fig. 9, we plot the test accuracy for these models.
@@
Comparing our results to DeltaProduct, we can see that the non-linearity introduced by the Fixed-Point dynamics allow for a slight improvement in the performance of the Householder products as the mixer components. Furthermore, we observe that the best performing mixer variant is still the Kroneckers model, which can successfully learn the state-tracking problem in all runs. Moreover, the FPMamba2 model demonsterates a better length generalization ability compared to FPMamba1, which we attribute to the improved underlying architecture used in the model. As shown in (Dao & Gu, 2024), Mamba2 has better recall capabilities, which can help with length generalization.
E.3 Sequential vs Parallel Fixed-Point Iteration
An important detail about the fixed-point framework proposed in this paper is that it is not convex. Therefore, the fixed-point is not necessarily unique, which can be problematic in autoregressive applications because there will be no guarantee that the parallel fixed-point during training will be the same as the sequential fixed-point used during inference (Schöne et al., 2025). In order to investigate this issue, we trained an FPMamba model on the $A_{5}$ task and compared the fixed-point computed sequentially and in parallel. We report the results in Fig. 10. We observe that the fixed-points are extremely similar, providing the possiblity of computing the fixed-point sequentially during inference.
@@ E.4 Modular Arithmetic Task Results
In Tab. 2, we evaluate modular arithmetic with brackets. We observe that a 2-layer FP-Mamba outperforms the baselines reported in (Grazzi et al., 2024) with a comparable number of parameters. In Fig. 11, we plot the validation accuracy as a function of the number of fixed-point iterations. We observe that the accuracy plateaus at $20$ iterations, which is significantly less than the shortest and longest sequence in the validation set. Therefore, the number of iterations required by FP-Mamba to reach its fixed point clearly does not scale with the sequence length in this task.
@@
2L FPMamba ($r=4$) 0.384
Table 2: The accuracy of various models on modular arithmetic with brackets. We adopt the reported numbers in (Grazzi et al., 2024) evaluating baselines the extended $[-1,1]$ eigenvalue range. Scores are commonly used scaled accuracies between $1.0$ and $0.0$ (random guessing). Highlighted is the best performance in each category.
Figure 11: Number of fixed-point iterations on the modular arithmetic task at test time. We report the validation accuracy after convergence for the number of fixed-point iterations caped at various values ranging from $2$ to $512$. The pink dashed line denotes the maximum sequence length during validation.
@@ Appendix F Low-Rank Expressiveness
where $\mathbb{W}_{d_{\omega}}$ is the set of words in the alphabet $[[d_{\omega}]]:=\{1,\dots,d_{\omega}\}~{}$ (i.e. $\mathbb{W}_{d_{\omega}}=\bigcup_{n\geq 0}[[d_{\omega}]]^{n}$ ) and for a given word $I=i_{1}\dots i_{n}$ with $S^{I}(\omega)_{[0,t]}$ we refer to the $I$th component of the signature tensor $S(\omega)_{[0,t]}$ i.e.
@@
[…]
@@
  • the initial value has independent standard Gaussian entries $[Y_{0}]_{\alpha}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1)$,
  • the weight matrices are distributed as $A^{i}\overset{\mathrm{iid}}{\sim}\frac{1}{\sqrt{Nr_{N}}}WM^{\top}$ with $W$ and $M$ independent $N\times r_{N}$ matrices having entries $[W]_{\alpha,\beta},[M]_{\alpha,\beta}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1)$,
@@
In our setting $\frac{1}{N}\langle A_{I}Y_{0},A_{J}Y_{0}\rangle_{\mathbb{R}^{N}}$ corresponds to a product graph $G_{I,J}$ corresponding to a ladder having $2|I|+2|J|$ edges as shown in figure 12. We can then use (Cirone et al., 2024a)[Prop. 2] to compute the square of the $L^{2}$ norm in equation Eq. 21, the only difference from the dense case is that half of the vertices (excluding the "middle" one) correspond to a space of dimension $r_{N}$ while the rest to the standard $N$.
Since $r_{N}\to\infty$ and given the scaling $N^{-1}\bigl{(}Nr_{N}\bigr{)}^{-\frac{|I|+|J|}{2}}$, the admissible pairings of $G_{I,J}$ not of order $o(1)$ are only the leading ones. These correspond to product graphs with $\frac{|I|+|J|}{2}$ $r_{N}$-dimensional vertices and $\frac{|I|+|J|}{2}+1$ $N$-dimensional vertices. By the same reasoning as in the full-rank case, these are found to be just the identity pairings.
@@
Since the total number of admissible pairings of $G_{I,J}\sqcup G_{I,J}$ is $\bigl{(}4(|I|+|J|)\bigr{)}!!$!, we conclude that equation 21 holds with $\kappa=4$ and $o(1):=\mathcal{O}\bigl{(}\frac{1}{\sqrt{N\wedge r_{N}}}\bigr{)}$.
@@
While the proof crucially uses the assumption $r_{N}\to\infty$ as $N\to\infty$, at the same time we have not provided an argument against $r_{N}$ not diverging. In figure 13 we present a counterexample, showing that if $r_{N}$ does not diverge then the asymptotics differ from the dense ones, in particular some symmetries are "lost", impossible to recover due to unavoidable noise.
Figure 13: Admissible pairing different from the "identity" paring, but still leading to maximal asymptotic scaling in the bounded $r_{N}$ case. Here, $I=12\neq 1112=J$, and we have highlighted in blue the vertices corresponding to the bounded dimension $r_{N}$. Recall that edges without arrows correspond to the matrix $\mathbf{I}$ (matrix of ones), and that two edges corresponding to matrices $A$ and $B$ which share direction and terminal vertices can be merged into the edge $A\odot B$.

### v3

@@ E.2 Long-Range State-Tracking
In this section, we investigate the ability of our proposed method in doing state tracking on longer sequences. Specifically, we will use the $A_{5}$ and $S_{5}$ datasets and train on sequence length 128, while evaluating for sequence lengths in the range $[2,512]$. We also implement the proposed Fixed-Point framework on Mamba2 (Dao & Gu, 2024), and we compared our method to DeltaProduct (Siems et al., 2025). In Fig. 9, we plot the test accuracy for these one-layer models.
@@
Comparing our results to DeltaProduct, we can see that the non-linearity introduced by the Fixed-Point dynamics allow for a slight improvement in the performance of the Householder products as the mixer components. Furthermore, we observe that the best performing mixer variant is still the Kroneckers model, which can successfully learn the state-tracking problem in all runs. Moreover, the FP-Mamba2 model demonsterates a better length generalization ability compared to FP-Mamba1, which we attribute to the improved underlying architecture used in the model. As shown in (Dao & Gu, 2024), Mamba2 has better recall capabilities, which can help with length generalization.
((a)) State Tracking on $A_{5}$ - FP-Mamba
((b)) State Tracking on $S_{5}$ - FP-Mamba
((c)) State Tracking on $A_{5}$ - FP-Mamba2
((d)) State Tracking on $S_{5}$ - FP-Mamba2
((e)) State Tracking on $A_{5}$ - DeltaProduct
((f)) State Tracking on $S_{5}$ - DeltaProduct
Figure 9: The state-tracking experiment for train sequence length $128$ and evaluation sequence length $[2,512]$. We omit the results of DPLR mixer due to poor performance. The figure presents (a, b) the results for FP-Mamba with Householder (H) and Kronecker (K) mixer, (c, d) the results for FP-Mamba2 with Householder (H) and Kronecker (K) mixer, and (e, f) the DeltaProduct method (Siems et al., 2025) for Householder mixers.
E.3 Reasoning on CatbAbI
In order to investigate the state-tracking ability of the fixed-point framework in a natural language setting, we perform experiments on the catbAbI dataset (Schlag et al., 2021b). catbAbI (concatenated-bAbI) is a reprocessing of the bAbI QA benchmark (Weston et al., 2015), where individual bAbI stories are stitched into one long, continuous sequence, so models must keep track of state across story boundaries. The task tries to stress-test the long-range state tracking and associative inference capabilities of sequence models beyond short, isolated contexts. Each sample in this dataset is a short story. At the end of each story, the model needs to choose a single word that is the answer to the question corresponding to the story. The responses include yes/no responses and the names of characters or locations in the story. We present the results in Table 3.
# Layers Mamba FP-Mamba-K FP-Mamba-H FP-Mamba-H FP-Mamba-H
($r=1$) ($r=2$) ($r=3$)
1 Layer 78.28% 79.93% 81.32% 81.60% 80.79%
2 Layers 87.08% 84.16% 89.08% 87.47% 89.55%
4 Layers 86.51%
Table 3: Test accuracy of the Mamba model vs. the FP-Mamba model for the Kronecker (K) and the Householder (H) channel mixers with $r\in\{1,2,3\}$ on the catbAbI dataset. We increase the number of layers to show the effect of having more layers on all models. The task benefits from the fixed-point dynamic, but increasing the number of layers seems to be suffering from diminishing returns.
In order to observe and compare the effect of more complex mixers with the number of layers, we use $1$, 22, and $4$ layers along with the Kronecker and Householder mixer with $r\in\{1,2,3\}$ reflections. Our investigation shows that increasing the number of layers seems to be reaching the point of diminishing returns very fast, while the fixed-point framework improves the performance. This observation seems to be in line with the findings of Saunshi et al. (2025), where the looped architecture seems to be providing a very helpful inductive bias for solving reasoning tasks. Comparing the performance of mixers, we observe that the Kronecker mixer under-performs compared to the Householder mixer, which we believe is in line with our observation in App. E.1, where the Kronecker mixer underperforms on tasks involving natural languages.
@@ E.4 Modular Arithmetic Task Results
Following Grazzi et al. (2024), we also evaluate FP-Mamba on the remaining unsolved task of the Chomsky Hierarchy of language problems introduced by Beck et al. (2024). Specifically, we focus on the mod arithmetic task with brackets. Following the setup of Grazzi et al. (2024), we train on sequence lengths $3$ to $40$ and report scaled accuracies on test sequences of lengths $40$ to $256$. For FP-Mamba, we use a $2$-layer model with $r=4$ reflections, i.e. the best performing model in the $A_{5}$ experiment.
In Tab. 4, we observe that a 2-layer FP-Mamba-H outperforms the baselines reported in (Grazzi et al., 2024) with a comparable number of parameters. In Fig. 10, we plot the validation accuracy as a function of the number of fixed-point iterations. We observe that the accuracy plateaus at $20$ iterations, which is significantly less than the shortest and longest sequence in the validation set. Therefore, the number of iterations required by FP-Mamba-H to reach its fixed point clearly does not scale with the sequence length in this task.
@@
2L FP-Mamba ($r=4$) 0.384
Table 4: The accuracy of various models on modular arithmetic with brackets. We adopt the reported numbers in (Grazzi et al., 2024) evaluating baselines the extended [1,1][-1,1] eigenvalue range. Scores are commonly used scaled accuracies between $1.0$ and $0.0$ (random guessing). Highlighted is the best performance in each category.
Figure 10: Number of fixed-point iterations on the modular arithmetic task at test time. We report the validation accuracy after convergence for the number of fixed-point iterations caped at various values ranging from 22 to $512$. The pink dashed line denotes the maximum sequence length during validation.
E.5 Effect of $\ell_{\text{max}}$ on test performance and number of iterations $\ell^{*}$
((a)) Effect of $\ell_{\text{max}}$
((b)) Number of Iterations $\ell^{*}$
Figure 11: The effect of $\ell_{\text{max}}$ on the performance of the model ((a)), and on the number of iterations $\ell^{*}$ ((b)) on the $A_{5}$ task. The vertical line denotes the train sequence length. All of the experiments are performed on FP-Mamba1 with a Householder mixer with $r=1$ reflections. Results are averaged across 4 runs.
In Fig. 11 we present the effect of the maximum number of iterations $\ell_{\text{max}}$ during training on the accuracy and the number of iterations $\ell^{*}$ during inference. As we observe, the general trend is that increasing $\ell_{\text{max}}$ improves the performance of the model. We attribute this observation to how well the model learns the task, as following Thm. 3.2, a condition for the gradients being a descent direction is for them to be computed at or close to the fixed-point. Consequently, we can see that when trained with a smaller number of iterations (small $\ell_{\text{max}}$), the model fails to fully utilize the fixed-point by adapting $\ell^{*}$ to the difficulty of the task.
E.6 Sequential vs Parallel Fixed-Point Iteration
An important detail about the fixed-point framework proposed in this paper is that it is not convex. Therefore, the fixed-point is not necessarily unique, which can be problematic in autoregressive applications because there are no guarantee that the parallel fixed-point during training will be the same as the sequential fixed-point used during inference (Schöne et al., 2025). In order to investigate this issue, we trained an FP-Mamba-H model on the $A_{5}$ task and compared the fixed-point computed sequentially and in parallel. We report the results in Fig. 12. We observe that the fixed-points are extremely similar, providing the possiblity of computing the fixed-point sequentially during inference.
@@ Appendix F Low-Rank Expressiveness
where $\mathbb{W}_{d_{\omega}}$ is the set of words in the alphabet $d_{\omega}]]:=\{1,\dots,d_{\omega}\}~$ (i.e. $\mathbb{W}_{d_{\omega}}=\bigcup_{n\geq 0}[[d_{\omega}]]^{n}$ ) and for a given word $I=i_{1}\dots i_{n}$ with $S^{I}(\omega)_{[0,t]}$ we refer to the $I$th component of the signature tensor $S(\omega)_{[0,t]}$ i.e.
@@
$\forall\epsilon>0,~\forall\alpha\in T(\mathbb{R}^{d}),~\exists N\geq 0,~\exists(v,(A^{i}),Y_{0})\in\mathbb{R}^{N}\times\mathcal{A}_{N}^{d}\times\mathcal{Y}_{N}~\text{s.t.}$
@@
  • the initial value has independent standard Gaussian entries $Y_{0}]_{\alpha}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1)$,
  • the weight matrices are distributed as $A^{i}\overset{\mathrm{iid}}{\sim}\frac{1}{\sqrt{Nr_{N}}}WM^{\top}$ with $W$ and $M$ independent $N\times r_{N}$ matrices having entries $W]_{\alpha,\beta},[M]_{\alpha,\beta}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1)$,
@@
In our setting $\frac{1}{N}\langle A_{I}Y_{0},A_{J}Y_{0}\rangle_{\mathbb{R}^{N}}$ corresponds to a product graph $G_{I,J}$ corresponding to a ladder having $2|I|+2|J|$ edges as shown in Fig. 13. We can then use (Cirone et al., 2024a)[Prop. 2] to compute the square of the $L^{2}$ norm in equation Eq. 22, the only difference from the dense case is that half of the vertices (excluding the "middle" one) correspond to a space of dimension $r_{N}$ while the rest to the standard $N$.
Since $r_{N}\to\infty$ and given the scaling $N^{-1}\bigl(Nr_{N}\bigr)^{-\frac{|I|+|J|}{2}}$, the admissible pairings of $G_{I,J}$ not of order $o(1)$ are only the leading ones. These correspond to product graphs with $\frac{|I|+|J|}{2}$ $r_{N}$-dimensional vertices and $\frac{|I|+|J|}{2}+1$ NN-dimensional vertices. By the same reasoning as in the full-rank case, these are found to be just the identity pairings.
@@
Since the total number of admissible pairings of $G_{I,J}\sqcup G_{I,J}$ is $\bigl(4(|I|+|J|)\bigr)!!$, we conclude that equation 22 holds with $\kappa=4$ and $o(1):=\mathcal{O}\bigl(\frac{1}{\sqrt{N\wedge r_{N}}}\bigr)$. ∎
@@
While the proof crucially uses the assumption $r_{N}\to\infty$ as $N\to\infty$, at the same time we have not provided an argument against $r_{N}$ not diverging. In Fig. 14 we present a counterexample, showing that if $r_{N}$ does not diverge then the asymptotics differ from the dense ones, in particular some symmetries are "lost", impossible to recover due to unavoidable noise.
Figure 14: Admissible pairing different from the "identity" paring, but still leading to maximal asymptotic scaling in the bounded $r_{N}$ case. Here, $I=12\neq 1112=J$, and we have highlighted in blue the vertices corresponding to the bounded dimension $r_{N}$. Recall that edges without arrows correspond to the matrix $\mathbf{I}$ (matrix of ones), and that two edges corresponding to matrices AA and $B$ which share direction and terminal vertices can be merged into the edge $A\odot B$.